An ideal transformer is supplied with voltage 240 V. It is found that the potential difference of the secondary coil is 12 V and the current is 3 A. Find the current in the primary coil.
1
Expert's answer
2020-05-11T20:21:15-0400

Ohm's low

V = I "\times" R

Data Given

V1 = 240 V

V2 = 12 V

I2 = 3A

I1 = ?

we know that , relationship of voltage and current,

Voltage directly proportional to current,

"V \propto I"


"\frac{V_{1}} {V_{2}}=\frac{I_{1}} {I_{2}}"


"I_{1} =I_{2}\times\frac{V_{1}} {V_{2}}"


"I_{1} =3 \times\frac{240} {12}"


"I_{1} = 60 A"